OpenStudy (anonymous):

Velocity is given by: v' = 1587 - 3v^2. How do I solve this equation and find the velocity as a function of time, leaving my answer implicitly? You can tell me which method I should use or leave an answer.

OpenStudy (anonymous):

It's a differential equation btw.

OpenStudy (anonymous):

you can rewrite that as dv/dt=1587-3v^2 then u have 1/dt=(1587-3v^2)/dv dt=dv/(1587-3v^2) now just integrate both sides